OpenStudy (anonymous):

can someone give me the formula to this question and explain how to do it please? "What is the arc measure of a sector with an area of 38.485 ft2 and in a circle with a radius of 7ft?"

OpenStudy (anonymous):

If the radius is 7, the entire circle would have area pi*7^2 = 49*pi =~ 153.938 But we only have 38.485, so divide that by (49*pi) and we have: 38.485 / (49*pi) =~ 0.25, that's almost exactly 1/4 of a circle (90 degrees or pi/2 radians)

OpenStudy (anonymous):

...idont get it.. @ninaesb

OpenStudy (anonymous):

thats what i found. do you have options ?

OpenStudy (anonymous):

no, i have to show work @ninaesb

OpenStudy (anonymous):

we have: 38.485 / (49*pi) =~ 0.25 do that .

OpenStudy (anonymous):

.25 is the answer? @ninaesb

OpenStudy (anonymous):

yeah , im guessing from that .

OpenStudy (anonymous):

@mikurout can u check and make sure?

OpenStudy (mikurout):

@shahlove Area of circle = pi*r^2 = 49pi for 2*pi*r of circumference we have area = 49pi so 153.93 which is 49pi = 2(pi)r so from this find what will be it for area of 38.485

OpenStudy (anonymous):

thats exaktly what i had , just in different words .

OpenStudy (mikurout):

@shahlove @ninaesb The answer is 10.99ft

OpenStudy (anonymous):

thanks @mikurout

OpenStudy (anonymous):

thanks @ninaesb

OpenStudy (anonymous):

no problem :)